Without solving the quadratic equation 5x^2 + 8x + 6 = 0, state whether or not this quadratic equation has real roots. Justify your answer.
Answer by stanbon(75887)

5x^2 + 8x + 6 = 0,
The determinant, b^2-4ac = 8^2-4(5)(6) is negative.
Therefore the equation has no Real roots.
Cheers,
Stan H.